# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

### v1

@@ Acknowledgements.
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.04(0.65, 0.35) -0.24(0.23, 0.77) -1.01(0.92, 0.08) -0.18(0.02, 0.98) -0.87(0.92, 0.08) 0.58(0.02, 0.98)
$2^{3}$ 9 -0.55(0.72, 0.28) -0.28(0.25, 0.75) -5.82(0.97, 0.03) -0.50(0.06, 0.94) -5.72(0.97, 0.03) -0.41(0.06, 0.94)
$2^{4}$ 10 -0.81(0.74, 0.26) 0.51(0.27, 0.73) 14.39(0.99, 0.01) 0.33(0.12, 0.88) 17.20(0.99, 0.01) 0.44(0.09, 0.91)
$2^{5}$ 11 -0.17(0.75, 0.25) 0.05(0.29, 0.71) 3.96(0.99, 0.01) -0.17(0.1, 0.90) -0.12(0.99, 0.01) 1.17(0.10, 0.90)
$2^{6}$ 12 0.30(0.75, 0.25) 0.54(0.30, 0.70) nan(0.99, 0.01) -1.16( 0.12, 0.88) -6.73(0.99, 0.01) 0.99(0.11, 0.89)
$2^{7}$ 13 nan(0.75, 0.25) 0.74(0.30,0.70) nan(0.99, 0.01) 0.30(0.12, 0.88) nan(0.99, 0.01) 2.20(0.02, 0.88)
$2^{8}$ 14 nan(0.75, 0.25) 0.40(0.30, 0.69) nan(0.99, 0.01) -7.03(0.12, 0.88) nan(0.99, 0.01) 3.54(0.02, 0.88)
$2^{9}$ 15 nan(0.75, 0.25) -0.31(0.31, 0.69) nan(0.99, 0.01) 2.88(0.12, 0.88) nan(0.99, 0.01) 1.96(0.02, 0.88)
$2^{10}$ 16 nan(0.75, 0.25) 0.02(0.31, 0.69) nan(0.99, 0.01) 5.28(0.12, 0.88) nan(0.99, 0.01) 4.93(0.02, 0.88)
$2^{11}$ 17 nan(0.75, 0.25) -0.30(0.31, 0.69) nan(0.99, 0.01) 2.52(0.12, 0.88) nan(0.99, 0.01) -2.36(0.02, 0.88)
$2^{12}$ 18 nan(0.75, 0.25) 0.48(0.31, 0.69) nan(0.99, 0.01) -3.50(0.12, 0.88) nan(0.99, 0.01) 0.81(0.02, 0.88)
$2^{13}$ 19 nan(0.75, 0.25) 0.38(0.31, 0.69) nan(0.99, 0.01) -9.08(0.12, 0.88) nan(0.99, 0.01) -2.22(0.02, 0.88)
$2^{14}$ 20 nan(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) -0.46(0.12, 0.88) nan(0.99, 0.01) 0.33(0.02, 0.88)
$2^{15}$ 21 nan(0.75, 0.25) 1.26(0.31, 0.69) nan(0.99, 0.01) 3.66(0.12, 0.88) nan(0.99, 0.01) 6.98(0.02, 0.88)
$2^{16}$ 22 nan(0.75, 0.25) -0.34(0.31, 0.69) nan(0.99, 0.01) 5.62(0.12, 0.88) nan(0.99, 0.01) -17.07(0.02, 0.88)
$2^{17}$ 23 nan(0.75, 0.25) 1.31(0.31, 0.69) nan(0.99, 0.01) 5.56(0.12, 0.88) nan(0.99, 0.01) 2.02(0.02, 0.88)
$2^{18}$ 24 nan(0.75, 0.25) 0.00(0.69, 0.31) nan(0.99, 0.01) 0.00(0.12, 0.88) nan(0.99, 0.01) 18.88(0.13, 0.87)
$2^{19}$ 25 nan(0.75, 0.25) 0.33(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) 6.42( 0.12, 0.88)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.27(0.31, 0.69) nan(0.99, 0.01) -6.81(0.12, 0.88) nan(0.99, 0.01) 6.75(0.12, 0.88)
$2^{2}$ 8 -0.57(0.64, 0.36) -0.01(0.23, 0.77) -0.87(0.92, 0.08) -0.02(0.02, 0.98) -0.88(0.92, 0.08) -0.01(0.02, 0.98)
$2^{3}$ 9 -0.61(0.72, 0.28) -0.16(0.22, 0.76) -5.82(0.97, 0.03) 0.04(0.06,0.94) -5.82(0.97, 0.03) 0.00(0.06, 0.94)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.05(0.27, 0.73) 14.66(0.99, 0.01) -0.26(0.09, 0.91) 14.95(0.99, 0.01) -0.27(0.09, 0.91)
$2^{5}$ 11 0.27(0.75, 0.25) -0.08(0.29, 0.71) 1.84(0.99, 0.01) -0.26(0.10, 0.90) 0.61(0.99, 0.01) 0.50(0.11, 0.89)
$2^{6}$ 12 0.04(0.75, 0.25) 0.02(0.3, 0.70) -0.72(0.99, 0.01) 1.32(0.11, 0.89) 1.68(0.99, 0.01) -0.27(0.12, 0.88)
$2^{7}$ 13 -0.07(0.75, 0.25) -0.07(0.31, 0.69) 1.35(0.99, 0.01) -0.24(0.12, 0.88) -4.87(0.99, 0.01) 0.40(0.12, 0.88)
$2^{8}$ 14 -0.13(0.75, 0.25) 0.12(0.31, 0.69) 0.09(0.99, 0.01) 0.43(0.12, 0.88) -4.04(0.99, 0.01) 1.91(0.12, 0.88)
$2^{9}$ 15 0.05(0.75, 0.25) 0.19(0.31, 0.69) -2.82(0.99, 0.01) -1.81(0.12, 0.88) nan(0.99, 0.01) 1.12(0.12, 0.88)
$2^{10}$ 16 -0.04(0.75, 0.25) -0.26(0.31, 0.69) nan(0.99, 0.01) 1.19(0.12, 0.88) nan(0.99, 0.01) -2.59(0.12, 0.88)
$2^{11}$ 17 0.18(0.75, 0.25) 0.04(0.31, 0.69) nan(0.99, 0.01) -1.68(0.12, 0.88) nan(0.99, 0.01) -0.85(0.12, 0.88)
$2^{12}$ 18 0.07(0.75, 0.25) -0.44(0.31, 0.69) nan(0.99, 0.01) 2.36(0.12, 0.88) nan(0.99, 0.01) -5.96(0.12, 0.88)
$2^{13}$ 19 0.12(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) 5.19(0.12, 0.88) nan(0.99, 0.01) 1.52(0.12, 0.88)
$2^{14}$ 20 -0.04(0.75, 0.25) 0.25(0.31, 0.69) nan(0.99, 0.01) 2.50(0.12, 0.88) nan(0.99, 0.01) -1.89(0.12, 0.88)
$2^{15}$ 21 0.06(0.75, 0.25) -0.33(0.31, 0.69) nan(0.99, 0.01) 0.41(0.12, 0.88) nan(0.99, 0.01) -3.65(0.12, 0.88)
$2^{16}$ 22 0.10(0.75, 0.25) -0.01(0.31, 0.69) nan(0.99, 0.01) -3.09(0.12, 0.88) nan(0.99, 0.01) -1.88(0.12, 0.88)
$2^{17}$ 23 -0.03(0.75, 0.25) -0.07(0.31, 0.69) nan(0.99, 0.01) 2.54(0.12, 0.88) nan(0.99, 0.01) -0.92(0.12, 0.88)
$2^{18}$ 24 0.11(0.75, 0.25) 0.07(0.31, 0.69) nan(0.99, 0.01) -0.47(0.88, 0.12) nan(0.99, 0.01) -0.46(0.88, 0.12)
$2^{19}$ 25 -0.08(0.75, 0.25) 0.09(0.31, 0.69) nan(0.99, 0.01) 1.81(0.12, 0.88) nan(0.99, 0.01) -0.89(0.12, 0.88)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.02(0.31, 0.69) nan(0.99, 0.01) -2.00(0.12, 0.88) nan(0.99, 0.01) 3.65(0.12, 0.88)
@@
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.44(0.65, 0.35) 0.43(0.21, 0.79) -1.01(0.92, 0.08) 0.06(0.01, 0.99) -0.87(0.92, 0.08) 0.31(0.02, 0.98)
$2^{3}$ 9 -0.55(0.72, 0.28) 0.29(0.27, 0.73) -5.82(0.97, 0.03) -0.59(0.02, 0.98) -5.72(0.97, 0.03) -0.92(0.02, 0.98)
$2^{4}$ 10 -0.81(0.74, 0.26) 2.47(0.32, 0.68) 14.39(0.99, 0.01) 2.59(0.02, 0.98) 17.20(0.99, 0.01) 1.55(0.02, 0.98)
$2^{5}$ 11 -0.17(0.75, 0.25) 0.67(0.35, 0.65) 3.96(0.99, 0.01) 1.63(0.03, 0.97) -0.12(0.99, 0.01) 1.89(0.03, 0.97)
$2^{6}$ 12 0.30(0.75, 0.25) -0.17(0.36, 0.64) -6.73(0.99, 0.01) 5.34(0.04, 0.96) -6.73(0.99, 0.01) -4.59(0.03, 0.97)
$2^{7}$ 13 nan(0.75, 0.25) -0.30(0.37, 0.63) nan(0.99, 0.01) 0.55(0.04, 0.96) nan(0.99, 0.01) 0.37(0.04, 0.96)
$2^{8}$ 14 nan(0.75, 0.25) -2.17(0.37, 0.63) nan(0.99, 0.01) 6.89(0.04, 0.96) nan(0.99, 0.01) 0.46(0.04, 0.96)
$2^{9}$ 15 nan(0.75, 0.25) 0.04(0.37, 0.63) nan(0.99, 0.01) -2.64(0.04, 0.96) nan(0.99, 0.01) 4.65(0.04, 0.96)
$2^{10}$ 16 nan(0.75, 0.25) 1.01(0.37, 0.63) nan(0.99, 0.01) 3.21(0.04, 0.96) nan(0.99, 0.01) -4.06(0.04, 0.96)
$2^{11}$ 17 nan(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) -5.46(0.04, 0.96) nan(0.99, 0.01) -0.93(0.04, 0.96)
$2^{12}$ 18 nan(0.75, 0.25) -0.55(0.37, 0.63) nan(0.99, 0.01) -4.80(0.04, 0.96) nan(0.99, 0.01) 9.42(0.04, 0.96)
$2^{13}$ 19 nan(0.75, 0.25) -0.17(0.37, 0.63) nan(0.99, 0.01) 2.18(0.04, 0.96) nan(0.99, 0.01) 7.83(0.04, 0.96)
$2^{14}$ 20 nan(0.75, 0.25) 0.56(0.37, 0.63) nan(0.99, 0.01) 5.32(0.04, 0.96) nan(0.99, 0.01) -0.86(0.04, 0.96)
$2^{15}$ 21 nan(0.75, 0.25) -0.11(0.37, 0.63) nan(0.99, 0.01) 5.89(0.04, 0.96) nan(0.99, 0.01) 3.92(0.04, 0.96)
$2^{16}$ 22 nan(0.75, 0.25) 0.24(0.37, 0.63) nan(0.99, 0.01) -12.26(0.04, 0.96) nan(0.99, 0.01) 9.60(0.04, 0.96)
$2^{17}$ 23 nan(0.75, 0.25) -0.94(0.37, 0.63) nan(0.99, 0.01) -2.94(0.04, 0.96) nan(0.99, 0.01) -5.61(0.04, 0.96)
$2^{18}$ 24 nan(0.75, 0.25) -1.64(0.26, 0.74) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) 5.92(0.04, 0.96)
$2^{19}$ 25 nan(0.75, 0.25) 1.47(0.27, 0.73) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) -4.92(0.04, 0.96)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -1.70(0.26, 0.74) nan(0.99, 0.01) -17.20(0.04, 0.96) nan(0.99, 0.01) 5.03(0.04, 0.96)
$2^{2}$ 8 -0.57(0.64, 0.36) -0.13(0.28, 0.72) -0.87(0.92, 0.08) 0.00(0.02, 0.98) -0.88(0.92, 0.08) 0.03(0.02, 0.98)
$2^{3}$ 9 -0.61(0.72, 0.28) 0.14(0.24, 0.76) -5.82(0.97, 0.03) -0.19(0.01, 0.99) -5.82(0.97, 0.03) -0.06(0.01, 0.99)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.03(0.25, 0.75) 14.66(0.99, 0.01) -0.22(0.02, 0.98) 14.95(0.99, 0.01) -0.02(0.02, 0.98)
$2^{5}$ 11 0.27(0.75, 0.25) 0.64(0.26, 0.74) 1.84(0.99, 0.01) 0.03(0.03, 0.97) 0.61(0.99, 0.01) 1.11(0.03, 0.97)
$2^{6}$ 12 0.04(0.75, 0.25) -0.05(0.26, 0.74) -0.72(0.99, 0.01) 1.21(0.04, 0.96) 1.68(0.99, 0.01) 0.68(0.03, 0.97)
$2^{7}$ 13 -0.07(0.75, 0.25) 0.17(0.26, 0.74) 1.35(0.99, 0.01) -1.22(0.04, 0.96) -4.87(0.99, 0.01) 0.24(0.03, 0.97)
$2^{8}$ 14 -0.13(0.75, 0.25) -0.17(0.26, 0.74) 0.09(0.99, 0.01) -3.19(0.04, 0.96) -4.04(0.99, 0.01) 0.32(0.04, 0.96)
$2^{9}$ 15 0.05(0.75, 0.25) -0.50(0.26, 0.74) 2.82(0.99, 0.01) -2.13(0.04, 0.96) nan(0.99, 0.01) -2.48(0.04, 0.96)
$2^{10}$ 16 -0.03(0.75, 0.25) -0.30(0.26, 0.74) nan(0.99, 0.01) -0.59(0.04, 0.96) nan(0.99, 0.01) -0.24(0.04, 0.96)
$2^{11}$ 17 0.18(0.75, 0.25) 0.47(0.26, 0.74) nan(0.99, 0.01) -2.93(0.04, 0.96) nan(0.99, 0.01) 0.05(0.04, 0.96)
$2^{12}$ 18 0.07(0.75, 0.25) 0.40(0.26, 0.74) nan(0.99, 0.01) 0.62(0.04, 0.96) nan(0.99, 0.01) -2.73(0.04, 0.96)
$2^{13}$ 19 0.11(0.75, 0.25) 0.14(0.26, 0.74) nan(0.99, 0.01) 0.12(0.04, 0.96) nan(0.99, 0.01) 6.03(0.04, 0.96)
$2^{14}$ 20 -0.03(0.75, 0.25) 0.62(0.26, 0.74) nan(0.99, 0.01) 2.52(0.04, 0.96) nan(0.99, 0.01) 0.28(0.04, 0.96)
$2^{15}$ 21 0.05(0.75, 0.25) -0.34(0.26, 0.74) nan(0.99, 0.01) 1.03(0.04, 0.96) nan(0.99, 0.01) -6.30(0.04, 0.96)
$2^{16}$ 22 0.10(0.75, 0.25) -0.59(0.26, 0.74) nan(0.99, 0.01) -3.21(0.04, 0.96) nan(0.99, 0.01) -2.40(0.04, 0.96)
$2^{17}$ 23 -0.03(0.75, 0.25) 0.49(0.26, 0.74) nan(0.99, 0.01) 0.45(0.04, 0.96) nan(0.99, 0.01) 2.72(0.04, 0.96)
$2^{18}$ 24 0.11(0.75, 0.25) 0.14(0.27, 0.73) nan(0.99, 0.01) -2.46(0.04, 0.96) nan(0.99, 0.01) -1.52(0.04, 0.96)
$2^{19}$ 25 0.08(0.75, 0.25) 0.01(0.27, 0.73) nan(0.99, 0.01) -1.95(0.04, 0.96) nan(0.99, 0.01) -0.72(0.04, 0.96)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.33(0.27, 0.73) nan(0.99, 0.01) -2.07(0.04, 0.96) nan(0.99, 0.01) -2.46(0.04, 0.96)
@@
Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{20}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$. Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots =$10^{6}$ is considered. The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails. The white cells denote those cases for which we did not carry a calculation, due to computational cost.

### v2

@@ Acknowledgements.
$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
# shots $\approx\kappa$ qubits PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1))
$2^{2}$ 8 -0.08(0.64, 0.36) -0.03(0.23, 0.77) 0.03(0.92, 0.08) -0.09(0.02, 0.98) 0.07(0.92, 0.08) -0.05(0.02, 0.98)
$2^{3}$ 9 0.01(0.72, 0.28) -0.17(0.24, 0.76) -0.05(0.97, 0.03) 0.12(0.06, 0.94) -0.02(0.97, 0.03) -0.22(0.06, 0.94)
$2^{4}$ 10 -0.02(0.74, 0.26) -0.26(0.27, 0.73) 0.31(0.99, 0.01) -0.48(0.12, 0.88) 0.48(0.99, 0.01) -0.17(0.09, 0.91)
$2^{5}$ 11 -0.06(0.75, 0.25) 0.07(0.29, 0.71) 1.85(0.99, 0.01) 0.70(0.10, 0.90) -0.10(0.99, 0.01) -0.68(0.10, 0.90)
$2^{6}$ 12 0.09(0.75, 0.25) 0.23(0.30, 0.70) nan(0.99, 0.01) -0.55(0.12, 0.88) nan(0.99, 0.01) 0.18(0.11, 0.89)
$2^{7}$ 13 -0.01(0.75, 0.25) 0.28(0.30, 0.70) nan(0.99, 0.01) 0.74(0.12, 0.88) nan(0.99, 0.01) 1.88(0.11, 0.88)
$2^{8}$ 14 0.15(0.75, 0.25) -0.09(0.3, 0.69) nan(0.99, 0.01) 2.95(0.12, 0.88) nan(0.99, 0.01) -1.04(0.12, 0.88)
$2^{9}$ 15 0.12(0.75, 0.25) 0.57(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) 2.59( 0.12, 0.88)
$2^{10}$ 16 -0.01(0.75, 0.25) -0.03(0.31, 0.69) nan(0.99, 0.01) -0.69(0.12, 0.88) nan(0.99, 0.01) 1.22(0.12, 0.88)
$2^{11}$ 17 -0.05(0.75, 0.25) -0.51(0.31, 0.69) nan(0.99, 0.01) 1.78(0.12, 0.88) nan(0.99, 0.01) -1.75(0.12, 0.88)
$2^{12}$ 18 0.02(0.75, 0.25) 0.06(0.31, 0.69) nan(0.99, 0.01) 4.37(0.12, 0.88) nan(0.99, 0.01) -0.00( 0.12, 0.88)
$2^{13}$ 19 0.10(0.75, 0.25) -0.38(0.31, 0.69) nan(0.99, 0.01) -2.56(0.12, 0.88) nan(0.99, 0.01) -0.03(0.12, 0.88)
$2^{14}$ 20 0.28(0.75, 0.25) 0.14(0.31, 0.69) nan(0.99, 0.01) 2.95(0.12, 0.88) nan(0.99, 0.01) -3.04(0.12, 0.88)
$2^{15}$ 21 0.21(0.75, 0.25) -0.26(0.31, 0.69) nan(0.99, 0.01) -0.37(0.12, 0.88) nan(0.99, 0.01) 3.27(0.12, 0.88)
$2^{16}$ 22 -0.11(0.75, 0.25) -0.24(0.31, 0.69) nan(0.99, 0.01) -1.35(0.12, 0.88) nan(0.99, 0.01) 2.36(0.12, 0.88)
$10^{5}$ $2^{17}$ 23 0.29(0.75, 0.25) -0.36(0.31, 0.69) nan(0.99, 0.01) 0.15(0.12, 0.88) nan(0.99, 0.01) 2.02(0.12, 0.88)
$2^{2}$ 8 -0.01(0.64, 0.36) -0.01(0.23, 0.77) 0.03(0.92, 0.08) 0.00(0.02, 0.98) 0.05(0.92, 0.08) -0.02(0.02, 0.98)
$2^{3}$ 9 0.00(0.72, 0.28) 0.07(0.24, 0.76) 0.02(0.97, 0.03) 0.01(0.06,0.94) -0.07(0.97, 0.03) 0.01(0.06, 0.94)
$2^{4}$ 10 0.00(0.74, 0.26) 0.02(0.27, 0.73) -0.07(0.99, 0.01) -0.06(0.09, 0.91) 0.12(0.99, 0.01) 0.14(0.09, 0.91)
$2^{5}$ 11 0.05(0.74, 0.26) -0.03(0.29, 0.71) 0.13(0.99, 0.01) 0.03(0.10, 0.90) 0.16(0.99, 0.01) 0.37(0.11, 0.89)
$2^{6}$ 12 0.07(0.75, 0.25) 0.01(0.30, 0.70) 0.99(0.99, 0.01) 0.68(0.11, 0.89) 0.03(0.99, 0.01) -0.03(0.12, 0.88)
$2^{7}$ 13 0.10(0.75, 0.25) -0.09(0.30, 0.70) 1.36(0.99, 0.01) -0.07(0.12, 0.88) 2.18(0.99, 0.01) 0.51(0.12, 0.88)
$2^{8}$ 14 -0.07(0.75, 0.25) 0.00(0.31, 0.69) 8.58(0.99, 0.01) -0.12(0.12, 0.88) 3.63(0.99, 0.01) -1.34(0.12, 0.88)
$2^{9}$ 15 0.05(0.75, 0.25) -0.08(0.31, 0.69) nan(0.99, 0.01) 0.23(0.12, 0.88) nan(0.99, 0.01) 0.25(0.12, 0.88)
$2^{10}$ 16 -0.14(0.75, 0.25) -0.12(0.31, 0.69) nan(0.99, 0.01) -0.31(0.12, 0.88) nan(0.99, 0.01) 0.71(0.12, 0.88)
$2^{11}$ 17 -0.09(0.75, 0.25) 0.03(0.31, 0.69) nan(0.99, 0.01) 0.88(0.12, 0.88) nan(0.99, 0.01) -0.31(0.12, 0.88)
$2^{12}$ 18 -0.04(0.75, 0.25) -0.11(0.31, 0.69) nan(0.99, 0.01) -0.32(0.12, 0.88) nan(0.99, 0.01) -0.82(0.12, 0.88)
$2^{13}$ 19 0.04(0.75, 0.25) 0.00(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) -0.86(0.12, 0.88)
$2^{14}$ 20 -0.12(0.75, 0.25) -0.07(0.31, 0.69) nan(0.99, 0.01) 0.84(0.12, 0.88) nan(0.99, 0.01) -0.06(0.12, 0.88)
$2^{15}$ 21 0.02(0.75, 0.25) 0.01(0.31, 0.69) nan(0.99, 0.01) 0.07(0.12, 0.88) nan(0.99, 0.01) 0.14(0.12, 0.88)
$2^{16}$ 22 0.09(0.75, 0.25) -0.05(0.31, 0.69) nan(0.99, 0.01) 0.75(0.12, 0.88) nan(0.99, 0.01) 0.21(0.12, 0.88)
$10^{6}$ $2^{17}$ 23 0.03(0.75, 0.25) 0.06(0.31, 0.69) nan(0.99, 0.01) 1.02(0.12, 0.88) nan(0.99, 0.01) 0.45(0.12, 0.88)
@@
$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
# shots $\approx\kappa$ qubits PFD(P(0), P(1)) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.08(0.64, 0.36) -0.08(0.28, 0.72) 0.04(0.92, 0.08) 0.27(0.02, 0.98) 0.07(0.92, 0.08) 0.26(0.02, 0.98)
$2^{3}$ 9 0.01(0.72, 0.28) -0.14(0.24, 0.76) -0.06(0.97, 0.03) 0.14(0.01, 0.99) -0.02(0.97, 0.03) 0.05(0.02, 0.98)
$2^{4}$ 10 -0.01(0.74, 0.26) 0.13(0.24, 0.76) 0.31(0.99, 0.01) 0.14(0.02, 0.98) 0.48(0.99, 0.01) -0.37(0.02, 0.98)
$2^{5}$ 11 -0.06(0.75, 0.25) -0.41(0.26, 0.74) 1.85(0.99, 0.01) 0.27(0.03, 0.97) -0.10(0.99, 0.01) -0.33(0.03, 0.97)
$2^{6}$ 12 0.08(0.75, 0.25) -0.62(0.26, 0.74) nan(0.99, 0.01) -0.48(0.04, 0.96) nan(0.99, 0.01) -0.28(0.03, 0.97)
$2^{7}$ 13 -0.01(0.75, 0.25) 0.76(0.26, 0.74) nan(0.99, 0.01) -1.63(0.04, 0.96) nan(0.99, 0.01) 1.86(0.04, 0.96)
$2^{8}$ 14 0.15(0.75, 0.25) 0.18(0.26, 0.74) nan(0.99, 0.01) -1.02(0.04, 0.96) nan(0.99, 0.01) 1.42(0.04, 0.96)
$2^{9}$ 15 0.12(0.75, 0.25) 0.20(0.27, 0.73) nan(0.99, 0.01) 5.05(0.04, 0.96) nan(0.99, 0.01) 4.95(0.04, 0.96)
$2^{10}$ 16 -0.01(0.75, 0.25) 0.25(0.26, 0.74) nan(0.99, 0.01) 2.12(0.04, 0.96) nan(0.99, 0.01) -3.28(0.04, 0.96)
$2^{11}$ 17 -0.05(0.75, 0.25) 0.85(0.26, 0.74) nan(0.99, 0.01) 0.70(0.04, 0.96) nan(0.99, 0.01) 0.53(0.04, 0.96)
$2^{12}$ 18 0.02(0.75, 0.25) -0.29(0.26, 0.74) nan(0.99, 0.01) 6.62(0.04, 0.96) nan(0.99, 0.01) 2.93(0.04, 0.96)
$2^{13}$ 19 0.10(0.75, 0.25) 1.14(0.27, 0.73) nan(0.99, 0.01) 5.23(0.04, 0.96) nan(0.99, 0.01) 0.42(0.04, 0.96)
$2^{14}$ 20 0.27(0.75, 0.25) -0.46(0.26, 0.74) nan(0.99, 0.01) 6.99(0.04, 0.96) nan(0.99, 0.01) 2.82(0.04, 0.96)
$2^{15}$ 21 0.21(0.75, 0.25) 0.50(0.26, 0.74) nan(0.99, 0.01) 1.87(0.04, 0.96) nan(0.99, 0.01) 1.49(0.04, 0.96)
$2^{16}$ 22 -0.11(0.75, 0.25) -0.46(0.27, 0.73) nan(0.99, 0.01) -1.35(0.04, 0.96) nan(0.99, 0.01) 0.76(0.04, 0.96)
$10^{5}$ $2^{17}$ 23 0.30(0.75, 0.25) -0.36(0.26, 0.74) nan(0.99, 0.01) 0.16(0.04, 0.96) nan(0.99, 0.01) 1.32(0.04, 0.96)
$2^{2}$ 8 -0.01(0.64, 0.36) 0.08(0.28, 0.72) 0.03(0.92, 0.08) 0.00(0.02, 0.98) 0.05(0.92, 0.08) -0.02(0.02, 0.98)
$2^{3}$ 9 0.00(0.72, 0.28) -0.21(0.24, 0.76) 0.02(0.97, 0.03) -0.02(0.01, 0.99) -0.07(0.97, 0.03) -0.11(0.01, 0.99)
$2^{4}$ 10 0.00(0.74, 0.26) -0.20(0.25, 0.75) -0.07(0.99, 0.01) -0.01(0.02, 0.98) 0.12(0.99, 0.01) 0.05(0.02, 0.98)
$2^{5}$ 11 0.06(0.75, 0.25) 0.19(0.26, 0.74) 0.12(0.99, 0.01) 0.49(0.03, 0.97) 0.16(0.99, 0.01) 0.17(0.03, 0.97)
$2^{6}$ 12 0.07(0.75, 0.25) -0.40(0.26, 0.74) 0.99(0.99, 0.01) 0.25 (0.04, 0.96) 0.03(0.99, 0.01) 0.23(0.03, 0.97)
$2^{7}$ 13 0.10(0.75, 0.25) 0.01(0.26, 0.74) 1.36(0.99, 0.01) 0.18(0.04, 0.96) 2.18(0.99, 0.01) 0.18(0.03, 0.97)
$2^{8}$ 14 -0.07(0.75, 0.25) -0.08(0.26, 0.74) 8.58(0.99, 0.01) -0.09(0.04, 0.96) 3.63(0.99, 0.01) -0.61(0.04, 0.96)
$2^{9}$ 15 0.05(0.75, 0.25) 0.24(0.26, 0.74) nan(0.99, 0.01) 0.84(0.04, 0.96) nan(0.99, 0.01) 0.01(0.04, 0.96)
$2^{10}$ 16 -0.14(0.75, 0.25) -0.21(0.26, 0.74) nan(0.99, 0.01) -0.93(0.04, 0.96) nan(0.99, 0.01) 0.63(0.04, 0.96)
$2^{11}$ 17 -0.09(0.75, 0.25) 0.00(0.26, 0.74) nan(0.99, 0.01) 1.95(0.04, 0.96) nan(0.99, 0.01) -0.37(0.04, 0.96)
$2^{12}$ 18 -0.04(0.75, 0.25) -0.14(0.26, 0.74) nan(0.99, 0.01) 0.54(0.04, 0.96) nan(0.99, 0.01) 2.01(0.04, 0.96)
$2^{13}$ 19 0.04(0.75, 0.25) 0.15(0.26, 0.74) nan(0.99, 0.01) -0.75(0.04, 0.96) nan(0.99, 0.01) -1.04(0.04, 0.96)
$2^{14}$ 20 -0.12(0.75, 0.25) 0.13(0.26, 0.74) nan(0.99, 0.01) 0.32(0.04, 0.96) nan(0.99, 0.01) 0.21(0.04, 0.96)
$2^{15}$ 21 0.02(0.75, 0.25) 0.01(0.26, 0.74) nan(0.99, 0.01) -0.66(0.04, 0.96) nan(0.99, 0.01) -0.37(0.04, 0.96)
$2^{16}$ 22 0.08(0.75, 0.25) 0.13(0.26, 0.74) nan(0.99, 0.01) 0.65(0.04, 0.96) nan(0.99, 0.01) 0.45(0.04, 0.96)
$10^{6}$ $2^{17}$ 23 0.02(0.75, 0.25) -0.24(0.26, 0.74) nan(0.99, 0.01) -2.53(0.04, 0.96) nan(0.99, 0.01) 0.52(0.04, 0.96)
@@
Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{17}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$. Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots = $10^{6}$ is considered. The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails. The white cells denote those cases for which we did not carry a calculation, due to computational cost.